### v1

# Toughness and $A_{\alpha}$-spectral radius in graphs thanks: This work was supported by the Natural Science Foundation of Shandong Province, China (ZR2023MA078)

Abstract
Let $\alpha\in[0,1)$, and let $G$ be a connected graph of order $n$ with $n\geq f(\alpha)$, where $f(\alpha)=6$ for $\alpha\in[0,\frac{2}{3}]$ and $f(\alpha)=\frac{4}{1-\alpha}$ for $\alpha\in(\frac{2}{3},1)$. A graph $G$ is said to be $t$-tough if $|S|\geq tc(G-S)$ for each subset $S$ of $V(G)$ with $c(G-S)\geq 2$, where $c(G-S)$ is the number of connected components in $G-S$. The $A_{\alpha}$-spectral radius of $G$ is denoted by $\rho_{\alpha}(G)$. In this paper, it is verified that $G$ is a 1-tough graph unless $G=K_{1}\vee(K_{n-2}\cup K_{1})$ if $\rho_{\alpha}(G)\geq\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$, where $\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$ equals the largest root of $x^{3}-((\alpha+1)n+\alpha-3)x^{2}+(\alpha n^{2}+(\alpha^{2}-\alpha-1)n-2\alpha+1)x-\alpha^{2}n^{2}+(3\alpha^{2}-\alpha+1)n-4\alpha^{2}+5\alpha-3=0$. Further, we present an $A_{\alpha}$-spectral radius condition for a graph to be a $t$-tough graph.
Keywords: graph; toughness; $A_{\alpha}$-spectral radius.
(2020) Mathematics Subject Classification: 05C50
1 Introduction
Let $G$ be an undirected simple graph with vertex set $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$ and edge set $E(G)$. The order of $G$ is the number $n=|V(G)|$ of its vertices and its size is the number $m=e(G)$ of its edges. A graph $G$ is called trivial if $|V(G)|=1$. For arbitrary $v\in V(G)$, the degree of $v$ in $G$ is defined as the number of edges which are adjacent to $v$ and denoted by $d_{G}(v)$. For a given subset $S$ of $V(G)$, we use $G[S]$ to denote the subgraph of $G$ induced by $S$, and write $G-S$ for $G[V(G)\setminus S]$. The complement of a graph $G$ is a graph $\overline{G}$ with the same vertex set as $G$, in which any two distinct vertices are adjacent if and only if they are nonadjacent in $G$. Let $G_{1}$ and $G_{2}$ be two vertex-disjoint graphs. We use $G_{1}\cup G_{2}$ to denote the disjoint union of $G_{1}$ and $G_{2}$. The join $G_{1}\vee G_{2}$ is the graph formed from $G_{1}\cup G_{2}$ by adding all possible edges between $V(G_{1})$ and $V(G_{2})$. We denote by $K_{n}$ a complete graph of order $n$.
Let $t$ be a nonnegative real number. A graph $G$ is said to be $t$-tough if $|S|\geq tc(G-S)$ for each subset $S$ of $V(G)$ with $c(G-S)\geq 2$, where $c(G-S)$ is the number of connected components in $G-S$. The toughness $t(G)$ of $G$ is the largest real number $t$ for which $G$ is $t$-tough, or is $\infty$ if $G$ is complete. This concept was first introduced by Chvátal [1] in 1973. Some results on toughness of graphs can be found in [2, 3, 4, 5, 6].
Given a graph $G$ of order $n$, the adjacency matrix $A(G)$ of $G$ is the $n\times n$ matrix in which entry $a_{ij}$ is 1 or 0 according to whether $v_{i}$ and $v_{j}$ are adjacent or not. The eigenvalues of $G$ are the eigenvalues of its adjacency matrix $A(G)$. Let $\lambda_{1}(G)\geq\lambda_{2}(G)\geq\dots\geq\lambda_{n}(G)$ be its eigenvalues in nonincreasing order. Note that the adjacency spectral radius of $G$ is equal to $\lambda_{1}(G)$, written as $\rho(G)$.
Let $D(G)$ denote the diagonal matrix of vertex degree of $G$. Then $L(G)=D(G)-A(G)$ and $Q(G)=D(G)+A(G)$ are called the Laplacian matrix and signless Laplacian matrix of $G$, respectively. For any $\alpha\in[0,1]$, Nikiforov [7] introduced the $A_{\alpha}$-matrix of $G$ as
$A_{\alpha}(G)=\alpha D(G)+(1-\alpha)A(G).$
It is easy to see that $A_{\alpha}(G)=A(G)$ if $\alpha=0$, and $A_{\alpha}(G)=\frac{1}{2}Q(G)$ if $\alpha=\frac{1}{2}$. Note that $A_{\alpha}(G)$ is a real symmetric nonnegative matrix. Hence, the eigenvalues of $A_{\alpha}(G)$ are real, which can be indexed in nonincreasing order as $\lambda_{1}(A_{\alpha}(G))\geq\lambda_{2}(A_{\alpha}(G))\geq\cdots\geq\lambda_{n}(A_{\alpha}(G))$. The largest eigenvalue $\lambda_{1}(A_{\alpha}(G))$ is called the $A_{\alpha}$-spectral radius of $G$, and denoted by $\rho_{\alpha}(G)$. Namely, $\rho_{\alpha}(G)=\lambda_{1}(A_{\alpha}(G))$. For some interesting spectral properties of $A_{\alpha}(G)$, we refer the reader to [7, 8, 9, 10, 11, 12, 13, 14].
Many authors [15, 16, 17, 18, 19, 20] obtained some results on spectral radius in graphs. Brouwer [21] investigated the relationship between toughness and eigenvalues and claimed that for any connected $d$-regular graph $G$, $t(G)>\frac{d}{\lambda}-2$, and he further conjectured that the lower bound $\frac{d}{\lambda}-2$ can be improved to $\frac{d}{\lambda}-1$, where $\lambda=\max_{2\leq i\leq n}|\lambda_{i}|$. Later, Gu [22] strengthened Brouwer’s result and claimed the lower bound $\frac{d}{\lambda}-2$ can be improved to $\frac{d}{\lambda}-\sqrt{2}$. Very recently, Gu [23] confirmed Brouwer’s conjecture. Fan, Lin and Lu [24] presented two adjacency spectral radius conditions for the existence of a $t$-tough graph. It is natural and interesting to find a sufficient condition for a graph to be $t$-tough in view of its $A_{\alpha}$-spectral radius. In this paper, we first show a $A_{\alpha}$-spectral radius condition to guarantee that a graph is 1-tough.
Theorem 1.1. Let $\alpha\in[0,1)$, and let $G$ be a connected graph of order $n$ with $n\geq f(\alpha)$, where
$f(\alpha)=\left\{\begin{array}[]{ll}6,&if\ \alpha\in[0,\frac{2}{3}];\\ \frac{4}{1-\alpha},&if\ \alpha\in(\frac{2}{3},1).\\ \end{array}\right.$
If $\rho_{\alpha}(G)\geq\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$, then $G$ is a 1-tough graph unless $G=K_{1}\vee(K_{n-2}\cup K_{1})$, where $\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$ equals the largest root of $x^{3}-((\alpha+1)n+\alpha-3)x^{2}+(\alpha n^{2}+(\alpha^{2}-\alpha-1)n-2\alpha+1)x-\alpha^{2}n^{2}+(3\alpha^{2}-\alpha+1)n-4\alpha^{2}+5\alpha-3=0$.
Further, it is natural and interesting to find an $A_{\alpha}$-spectral radius condition to ensure that a graph is a $t$-tough graph. Next, we present an $A_{\alpha}$-spectral radius condition for a graph to be $t$-tough.
Theorem 1.2. Let $\alpha\in[\frac{1}{2},\frac{3}{4})$, and let $t$ be a positive integer. If $G$ be a connected graph of order $n\geq\max\{4t^{2}+6t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}$ with $\rho_{\alpha}(G)\geq\rho_{\alpha}(K_{2t-1}\vee(K_{n-2t}\cup K_{1}))$, then $G$ is a $t$-tough graph unless $G=K_{2t-1}\vee(K_{n-2t}\cup K_{1})$.
2 Preliminaries
In this section, we put forward some necessary lemmas, which play a key role in proving our main results.
Lemma 2.1 (Nikiforov [7]). Let $K_{n}$ be a complete graph of order $n$. Then
$\rho_{\alpha}(K_{n})=n-1.$
Lemma 2.2 (Nikiforov [7]). If $G$ is a connected graph, and $H$ is a proper subgraph of $G$, then
$\rho_{\alpha}(G)>\rho_{\alpha}(H).$
Lemma 2.3 (Zhao, Huang and Wang [14]). Let $\alpha\in[0,1)$, and let $n_{1}\geq n_{2}\geq\cdots\geq n_{t}$ be positive integers with $n=\sum_{i=1}^{t}n_{i}+s$ and $n_{1}\leq n-s-t+1$. Then
$\rho_{\alpha}(K_{s}\vee(K_{n_{1}}\cup K_{n_{2}}\cup\cdots\cup K_{n_{t}}))\leq\rho_{\alpha}(K_{s}\vee(K_{n-s-t+1}\cup(t-1)K_{1})),$
where the equality holds if and only if $(n_{1},n_{2},\ldots,n_{t})=(n-s-t+1,1,\ldots,1)$.
Let $M$ be a real matrix whose rows and columns are indexed by $V=\{1,2,\ldots,n\}$. Assume that $M$, with respect to the partition $\pi:V=V_{1}\cup V_{2}\cup\cdots\cup V_{t}$, can be written as
$M=\left(\begin{array}[]{cccc}M_{11}&M_{12}&\cdots&M_{1t}\\ M_{21}&M_{22}&\cdots&M_{2t}\\ \vdots&\vdots&\ddots&\vdots\\ M_{t1}&M_{t2}&\cdots&M_{tt}\\ \end{array}\right),$
where $M_{ij}$ denotes the submatrix (block) of $M$ formed by rows in $V_{i}$ and columns in $V_{j}$. Let $b_{ij}$ denote the average row sum of $M_{ij}$, namely, $b_{ij}$ is the sum of all entries in $M_{ij}$ divided by the number of rows. Then matrix $M_{\pi}=(b_{ij})$ is called the quotient matrix of $M$. If the row sum of every block $M_{ij}$ is a constant, then the partition is equitable.
Lemma 2.4 (Haemers [25], You, Yang, So and Xi [26]). Let $M$ be a real matrix with an equitable partition $\pi$, and let $M_{\pi}$ be the corresponding quotient matrix. Then each eigenvalue of $M_{\pi}$ is an eigenvalue of $M$. Furthermore, if $M$ is nonnegative, then the spectral radius of $M_{\pi}$ equals the spectral radius of $M$.
Lemma 2.5 (Alhevaz, Baghipur, Ganie and Das [27]). Let $G$ be a graph of order $n$ with $m$ edges, with no isolated vertices and let $\alpha\in[\frac{1}{2},1]$. Then
$\rho_{\alpha}(G)\leq\frac{2m(1-\alpha)}{n-1}+\alpha n-1.$
If $\alpha\in(\frac{1}{2},1)$ and $G$ is connected, equality holds if and only if $G=K_{n}$.
3 The proof of Theorem 1.1
In this section, we verify Theorem 1.1, which establishes an $A_{\alpha}$-spectral radius condition for a graph to be 1-tough.
Proof of Theorem 1.1. Suppose to the contrary that $G$ is not a 1-tough graph, there exists some nonempty subset $S\subseteq V(G)$ such that $c(G-S)\geq|S|+1$. Let $|S|=s$. Then $G$ is a spanning subgraph of $G_{s}^{1}=K_{s}\vee(K_{n_{1}}\cup K_{n_{2}}\cup\cdots\cup K_{n_{s+1}})$ for some integers $n_{1}\geq n_{2}\geq\cdots\geq n_{s+1}$ with $\sum\limits_{i=1}^{s+1}n_{i}=n-s$. According to Lemma 2.2, we have
$\rho_{\alpha}(G)\leq\rho_{\alpha}(G_{s}^{1})$ (3.1)
with equality if and only if $G=G_{s}^{1}$.
Let $G_{s}^{2}=K_{s}\vee(K_{n-2s}\cup sK_{1})$. Using Lemma 2.3, we get
$\rho_{\alpha}(G_{s}^{1})\leq\rho_{\alpha}(G_{s}^{2}),$ (3.2)
where the equality holds if and only if $(n_{1},n_{2},\ldots,n_{s+1})=(n-2s,1,\ldots,1)$. Let $\varphi(x)=x^{3}-((\alpha+1)n+\alpha-3)x^{2}+(\alpha n^{2}+(\alpha^{2}-\alpha-1)n-2\alpha+1)x-\alpha^{2}n^{2}+(3\alpha^{2}-\alpha+1)n-4\alpha^{2}+5\alpha-3$. Then $\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$ equals the largest root of $\varphi(x)=0$. We are to verify the following claim.
Claim 1. If $\alpha\in[0,1)$ and $n\geq f(\alpha)$, then we obtain
$\rho_{\alpha}(G_{s}^{2})\leq\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$
with equality if and only if $G_{s}^{2}=K_{1}\vee(K_{n-2}\cup K_{1})$.
Proof. Recall that $G_{s}^{2}=K_{s}\vee(K_{n-2s}\cup sK_{1})$. Obviously, $1\leq s\leq\frac{n-1}{2}$.
Consider the partition $V(G_{s}^{2})=V(sK_{1})\cup V(K_{n-2s})\cup V(K_{s})$. The corresponding quotient matrix of $A_{\alpha}(G_{s}^{2})$ is equal to
$B_{1}=\left(\begin{array}[]{ccc}\alpha s&0&(1-\alpha)s\\ 0&n+(\alpha-2)s-1&(1-\alpha)s\\ (1-\alpha)s&(1-\alpha)(n-2s)&\alpha n-\alpha s+s-1\\ \end{array}\right).$
Then the characteristic polynomial of $B_{1}$ is
$\varphi_{B_{1}}(x)=$ $x^{3}-((\alpha+1)n+(\alpha-1)s-2)x^{2}$
$+(\alpha n^{2}+(\alpha^{2}s-\alpha-1)n-s^{2}-(2\alpha-1)s+1)x$
$-\alpha^{2}sn^{2}+(2\alpha^{2}-2\alpha+1)s^{2}n+(\alpha^{2}+\alpha)sn$
$-(3\alpha^{2}-5\alpha+2)s^{3}-(\alpha^{2}-\alpha+1)s^{2}-\alpha s.$ (3.3)
Since the partition $V(G_{s}^{2})=V(sK_{1})\cup V(K_{n-2s})\cup V(K_{s})$ is equitable, it follows from Lemma 2.4 that $\rho_{\alpha}(G_{s}^{2})$ is the largest root of $\varphi_{B_{1}}(x)=0$. Namely, $\varphi_{B_{1}}(\rho_{\alpha}(G_{s}^{2}))=0$. Let $\eta_{1}=\rho_{\alpha}(G_{s}^{2})\geq\eta_{2}\geq\eta_{3}$ be the three roots of $\varphi_{B_{1}}(x)=0$ and $Q=\mbox{diag}(s,n-2s,s)$. One checks that
$Q^{\frac{1}{2}}B_{1}Q^{-\frac{1}{2}}=\left(\begin{array}[]{ccc}\alpha s&0&(1-\alpha)s\\ 0&n+(\alpha-2)s-1&(1-\alpha)s^{\frac{1}{2}}(n-2s)^{\frac{1}{2}}\\ (1-\alpha)s&(1-\alpha)s^{\frac{1}{2}}(n-2s)^{\frac{1}{2}}&\alpha n-\alpha s+s-1\\ \end{array}\right)$
is symmetric, and also contains
$\left(\begin{array}[]{ccc}\alpha s&0\\ 0&n+(\alpha-2)s-1\\ \end{array}\right)$
as its submatrix. Since $Q^{\frac{1}{2}}B_{1}Q^{-\frac{1}{2}}$ and $B_{1}$ have the same eigenvalues, the Cauchy interlacing theorem (cf. [25]) implies that
$\eta_{2}\leq n+(\alpha-2)s-1<n-2.$ (3.4)
If $s=1$, then $G_{s}^{2}=K_{1}\vee(K_{n-2}\cup K_{1})$ and $\varphi_{B_{1}}(x)=\varphi(x)$. And so $\rho_{\alpha}(G_{s}^{2})=\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$. In what follows, we always assume $s\geq 2$.
Note that $K_{n-1}$ is a proper subgraph of $K_{1}\vee(K_{n-2}\cup K_{1})$. In terms of (3.4), Lemmas 2.1 and 2.2, we get
$\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))>\rho_{\alpha}(K_{n-1})=n-2>\eta_{2}.$ (3.5)
Write $\theta(n)=\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$. Note that $\varphi(\theta(n))=0$. A simple calculation yields that
$\varphi_{B_{1}}(\theta(n))=$ $\varphi_{B_{1}}(\theta(n))-\varphi(\theta(n))$
$=$ $(s-1)((1-\alpha)(\theta(n))^{2}+(\alpha^{2}n-2\alpha-s)\theta(n)-\alpha^{2}n^{2}$
$+(2\alpha^{2}-2\alpha+1)sn+(3\alpha^{2}-\alpha+1)n-(3\alpha^{2}-5\alpha+2)s^{2}$
$-(4\alpha^{2}-6\alpha+3)s-4\alpha^{2}+5\alpha-3).$ (3.6)
The following proof will be divided into two cases.
Case 1. $0\leq\alpha\leq\frac{2}{3}$.
Let $h(x)=(1-\alpha)x^{2}+(\alpha^{2}n-2\alpha-s)x-\alpha^{2}n^{2}+(2\alpha^{2}-2\alpha+1)sn+(3\alpha^{2}-\alpha+1)n-(3\alpha^{2}-5\alpha+2)s^{2}-(4\alpha^{2}-6\alpha+3)s-4\alpha^{2}+5\alpha-3$ be a real function in $x$. Then the symmetry axis of $h(x)$ is $x=\frac{-\alpha^{2}n+2\alpha+s}{2(1-\alpha)}$. Obviously, $h(x)$ is increasing when $x\geq\frac{-\alpha^{2}n+2\alpha+s}{2(1-\alpha)}$. By plugging the value $\theta(n)$ into $x$ of $h(x)$, we obtain
$h(\theta(n))=$ $(1-\alpha)(\theta(n))^{2}+(\alpha^{2}n-2\alpha-s)\theta(n)-\alpha^{2}n^{2}$
$+(2\alpha^{2}-2\alpha+1)sn+(3\alpha^{2}-\alpha+1)n-(3\alpha^{2}-5\alpha+2)s^{2}$
$-(4\alpha^{2}-6\alpha+3)s-4\alpha^{2}+5\alpha-3$
and it follows from (3) that
$\varphi_{B_{1}}(\theta(n))=(s-1)h(\theta(n)).$ (3.7)
Note that $n-2s\geq 1$. We have $n\geq 2s+1$. Together with $s\geq 2$ and (3.5), we conclude
$\frac{-\alpha^{2}n+2\alpha+s}{2(1-\alpha)}<n-2<\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))=\theta(n),$
and so
$h(\theta(n))>$ $h(n-2)$
$=$ $(1-\alpha)(n-2)^{2}+(\alpha^{2}n-2\alpha-s)(n-2)-\alpha^{2}n^{2}$
$+(2\alpha^{2}-2\alpha+1)sn+(3\alpha^{2}-\alpha+1)n-(3\alpha^{2}-5\alpha+2)s^{2}$
$-(4\alpha^{2}-6\alpha+3)s-4\alpha^{2}+5\alpha-3$
$=$ $(1-\alpha)n^{2}+((2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3)n-(3\alpha^{2}-5\alpha+2)s^{2}$
$-(4\alpha^{2}-6\alpha+1)s-4\alpha^{2}+5\alpha+1.$ (3.8)
Let $p(x,s)=(1-\alpha)x^{2}+((2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3)x-(3\alpha^{2}-5\alpha+2)s^{2}-(4\alpha^{2}-6\alpha+1)s-4\alpha^{2}+5\alpha+1$ be a real function in $x$. According to (3), we get
$h(\theta(n))>p(n,s).$ (3.9)
Clearly, $p(x,s)$ is increasing when $x\geq-\frac{(2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3}{2(1-\alpha)}$. If $s\geq 3$, then we can verify that
$-\frac{(2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3}{2(1-\alpha)}<2s+1\leq n,$
and so
$p(n,s)\geq$ $p(2s+1,s)$
$=$ $(1-\alpha)(2s+1)^{2}+((2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3)(2s+1)$
$-(3\alpha^{2}-5\alpha+2)s^{2}-(4\alpha^{2}-6\alpha+1)s-4\alpha^{2}+5\alpha+1$
$=$ $(s^{2}-3)\alpha^{2}-(3s^{2}-2s-5)\alpha+2s^{2}-3s-1$
$\geq$ $\frac{4}{9}(s^{2}-3)-\frac{2}{3}(3s^{2}-2s-5)+2s^{2}-3s-1$
$=$ $\frac{4s^{2}-15s+9}{9}$
$\geq$ $0,$
where the last two inequalities hold from $\frac{3s^{2}-2s-5}{2(s^{2}-3)}>\frac{2}{3}\geq\alpha$ and $s\geq 3$, respectively.
If $s=2$, then $-\frac{(2\alpha^{2}-2\alpha)s+\alpha^{2}+\alpha-3}{2(1-\alpha)}=-\frac{5\alpha^{2}-3\alpha-3}{2(1-\alpha)}<6=f(\alpha)\leq n$ due to $0\leq\alpha\leq\frac{2}{3}$. Hence, for $0\leq\alpha\leq\frac{2}{3}$ and $n\geq f(\alpha)=6$, we get
$p(n,2)\geq p(6,2)=6\alpha^{2}-17\alpha+9>0.$
Thus, we obtain $p(n,s)>0$ for $2\leq s\leq\frac{n-1}{2}$. Together with (3.7), (3.9) and $s\geq 2$, we have
$\varphi_{B_{1}}(\theta(n))=(s-1)h(\theta(n))>(s-1)p(n,s)\geq 0.$
Recall that $\rho_{\alpha}(G_{s}^{2})$ is the largest root of $\varphi_{B_{1}}(x)=0$. As $\eta_{2}<n-2<\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))=\theta(n)$ (see (3.5)), it follows from (3), $\frac{-\alpha^{2}n+2\alpha+s}{2(1-\alpha)}<n-2<\theta(n)$ and $\varphi_{B_{1}}(\theta(n))>0$ that
$\rho_{\alpha}(G_{s}^{2})<\theta(n)=\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$
for $2\leq s\leq\frac{n-1}{2}$.
Case 2. $\frac{2}{3}<\alpha<1$.
In terms of (3), we obtain
$\varphi_{B_{1}}(n-2)=$ $(n-2)^{3}-((\alpha+1)n+(\alpha-1)s-2)(n-2)^{2}$
$+(\alpha n^{2}+(\alpha^{2}s-\alpha-1)n-s^{2}-(2\alpha-1)s+1)(n-2)$
$-\alpha^{2}sn^{2}+(2\alpha^{2}-2\alpha+1)s^{2}n+(\alpha^{2}+\alpha)sn$
$-(3\alpha^{2}-5\alpha+2)s^{3}-(\alpha^{2}-\alpha+1)s^{2}-\alpha s$
$=$ $(3\alpha-2)(1-\alpha)s^{3}+((2\alpha^{2}-2\alpha)n-\alpha^{2}+\alpha+1)s^{2}$
$+((1-\alpha)n^{2}-(\alpha^{2}-3\alpha+3)n-\alpha+2)s$
$+(\alpha-1)n^{2}-(2\alpha-3)n-2$
$:=\Phi(s,n).$
Thus, we obtain
$\frac{\partial\Phi(s,n)}{\partial s}=$ $3(3\alpha-2)(1-\alpha)s^{2}+2((2\alpha^{2}-2\alpha)n-\alpha^{2}+\alpha+1)s$
$+(1-\alpha)n^{2}-(\alpha^{2}-3\alpha+3)n-\alpha+2.$
Note that $\frac{2}{3}<\alpha<1$ and $n\geq f(\alpha)=\frac{4}{1-\alpha}$. By a simple calculation, we have
$\frac{\partial\Phi(s,n)}{\partial s}\Big{|}_{s=2}=$ $(1-\alpha)n^{2}+(7\alpha^{2}-5\alpha-3)n-40\alpha^{2}+63\alpha-18$
$\geq$ $(1-\alpha)\left(\frac{4}{1-\alpha}\right)^{2}+(7\alpha^{2}-5\alpha-3)\left(\frac{4}{1-\alpha}\right)-40\alpha^{2}+63\alpha-18$
$=$ $\frac{1}{1-\alpha}(40\alpha^{3}-75\alpha^{2}+61\alpha-14)$
$>$ $0,$
and
$\frac{\partial\Phi(s,n)}{\partial s}\Big{|}_{s=\frac{n-1}{2}}=$ $\frac{1}{4}((1-\alpha)(\alpha-2)n^{2}+(2\alpha^{2}-6\alpha-4)n-5\alpha^{2}+7\alpha+6)$
$\leq$ $\frac{1}{4}\left((1-\alpha)(\alpha-2)\left(\frac{4}{1-\alpha}\right)^{2}+(2\alpha^{2}-6\alpha-4)\left(\frac{4}{1-\alpha}\right)-5\alpha^{2}+7\alpha+6\right)$
$=$ $\frac{1}{4(1-\alpha)}(5\alpha^{3}-4\alpha^{2}-7\alpha-42)$
$<$ $0$
This implies that $\varphi_{B_{1}}(n-2)=\Phi(s,n)\geq\min\left\{\Phi(2,n),\Phi\left(\frac{n-1}{2},n\right)\right\}$ because the leading coefficient of $\Phi(s,n)$ (view as a cubic polynomial of $s$) is positive, and $2\leq s\leq\frac{n-1}{2}$. According to $\frac{2}{3}<\alpha<1$ and $n\geq f(\alpha)=\frac{4}{1-\alpha}$, we obtain
$\Phi(2,n)=$ $(1-\alpha)n^{2}+(6\alpha^{2}-4\alpha-3)n-28\alpha^{2}+42\alpha-10$
$\geq$ $(1-\alpha)\left(\frac{4}{1-\alpha}\right)^{2}+(6\alpha^{2}-4\alpha-3)\left(\frac{4}{1-\alpha}\right)-28\alpha^{2}+42\alpha-10$
$=$ $\frac{2}{1-\alpha}(14\alpha^{3}-23\alpha^{2}+18\alpha-3)$
$>$ $0,$
and
$\Phi\left(\frac{n-1}{2},n\right)=$ $\frac{1}{8}((\alpha^{2}-3\alpha+2)n^{3}-(5\alpha^{2}-19\alpha+16)n^{2}$
$+(3\alpha^{2}-25\alpha+34)n+\alpha^{2}+\alpha-20)$
$\geq$ $\frac{1}{8}((\alpha^{2}-3\alpha+2)\left(\frac{4}{1-\alpha}\right)^{3}-(5\alpha^{2}-19\alpha+16)\left(\frac{4}{1-\alpha}\right)^{2}$
$+(3\alpha^{2}-25\alpha+34)\left(\frac{4}{1-\alpha}\right)+\alpha^{2}+\alpha-20)$
$=$ $\frac{1}{8(1-\alpha)^{2}}(\alpha^{4}-13\alpha^{3}+11\alpha^{2}+45\alpha-12)$
$>$ $0$
Consequently, we conclude $\varphi_{B_{1}}(n-2)\geq\min\left\{\Phi(2,n),\Phi\left(\frac{n-1}{2},n\right)\right\}>0$ for $2\leq s\leq\frac{n-1}{2}$. As $\eta_{2}<n-2<\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$ (see (3.5)), we have
$\rho_{\alpha}(G_{s}^{2})<n-2<\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$
for $2\leq s\leq\frac{n-1}{2}$. This verifies Claim 1. $\Box$
In terms of (3.1), (3.2) and Claim 1, we obtain
$\rho_{\alpha}(G)\leq\rho_{\alpha}(K_{1}\vee(K_{n-2}\cup K_{1}))$
with equality if and only if $G=K_{1}\vee(K_{n-2}\cup K_{1})$. Which is a contradiction to the $A_{\alpha}$-spectral condition of Theorem 1.1. This completes the proof of Theorem 1.1. $\Box$
4 The proof of Theorem 1.2
In this section, we prove Theorem 1.2, which provides an $A_{\alpha}$-spectral radius condition for a graph to be $t$-tough.
Proof of Theorem 1.2. Suppose to the contrary that $G$ is not a $t$-tough graph, there exists some nonempty subset $S\subseteq V(G)$ such that $tc(G-S)-1\geq|S|$. Let $|S|=s$ and $c(G-S)=c$. Then $G$ is a spanning subgraph of $G_{1}=K_{tc-1}\vee(K_{n_{1}}\cup K_{n_{2}}\cup\cdots\cup K_{n_{c}})$ for some integers $n_{1}\geq n_{2}\geq\cdots\geq n_{c}$ with $\sum\limits_{i=1}^{c}n_{i}=n-tc+1$. In terms of Lemma 2.2, we conclude
$\rho_{\alpha}(G)\leq\rho_{\alpha}(G_{1})$ (4.1)
with equality if and only if $G=G_{1}$.
Let $G_{2}=K_{tc-1}\vee(K_{n-(t+1)c+2}\cup(c-1)K_{1})$. By virtue of Lemma 2.3, we obtain
$\rho_{\alpha}(G_{1})\leq\rho_{\alpha}(G_{2}),$ (4.2)
where the equality holds if and only if $(n_{1},n_{2},\ldots,n_{c})=(n-(t+1)c+2,1,\ldots,1)$.
If $c=2$, then we have $G_{2}=K_{2t-1}\vee(K_{n-2t}\cup K_{1})$. According to (4.1) and (4.2), we get
$\rho_{\alpha}(G)\leq\rho_{\alpha}(K_{2t-1}\vee(K_{n-2t}\cup K_{1})),$
where the equality holds if and only if $G=K_{2t-1}\vee(K_{n-2t}\cup K_{1})$. In what follows, we consider $c\geq 3$.
Recall that $G_{2}=K_{tc-1}\vee(K_{n-(t+1)c+2}\cup(c-1)K_{1})$. According to $\alpha\in[\frac{1}{2},\frac{3}{4})$ and Lemma 2.5, we obtain
$\rho_{\alpha}(G_{2})\leq$ $\frac{2e(G_{2})(1-\alpha)}{n-1}+\alpha n-1$
$=$ $\frac{(1-\alpha)((n-c+1)(n-c)+2(tc-1)(c-1))}{n-1}+\alpha n-1$
$=$ $\frac{1-\alpha}{n-1}((2t+1)c^{2}-(2n+2t+3)c+n^{2}+n+2)+\alpha n-1.$ (4.3)
Let $\varphi(c)=(2t+1)c^{2}-(2n+2t+3)c+n^{2}+n+2$. Note that $n\geq(t+1)c-1$. We conclude $3\leq c\leq\frac{n+1}{t+1}$. By a direct computation, it follows from $n\geq\max\{4t^{2}+6t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}$ that
$\varphi(3)-\varphi\left(\frac{n+1}{t+1}\right)=\frac{(n-3t-2)(n-4t^{2}-6t-1)}{(t+1)^{2}}\geq 0,$
which implies that $\varphi(c)$ attains its maximum value at $c=3$ when $3\leq c\leq\frac{n+1}{t+1}$. Combining this with (4), $\alpha\in[\frac{1}{2},\frac{3}{4})$ and $n\geq\max\{4t^{2}+6t+1,\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}\}$, we get
$\rho_{\alpha}(G_{2})\leq$ $\frac{(1-\alpha)\varphi(3)}{n-1}+\alpha n-1$
$=$ $\frac{(1-\alpha)(n^{2}-5n+12t+2)}{n-1}+\alpha n-1$
$=$ $n-2+\frac{-(3-4\alpha)n+12t(1-\alpha)-2\alpha+1}{n-1}$
$\leq$ $n-2+\frac{-(3-4\alpha)\cdot\frac{12t(1-\alpha)-2\alpha+1}{3-4\alpha}+12t(1-\alpha)-2\alpha+1}{n-1}$
$=$ $n-2.$ (4.4)
Since $K_{n-1}$ is a proper subgraph of $K_{2t-1}\vee(K_{n-2t}\cup K_{1})$, we conclude
$\rho_{\alpha}(K_{2t-1}\vee(K_{n-2t}\cup K_{1}))>\rho_{\alpha}(K_{n-1})=n-2$ (4.5)
by Lemmas 2.1 and 2.2. It follows from (4.1), (4.2), (4) and (4.5) that
$\rho_{\alpha}(G)\leq\rho_{\alpha}(G_{1})\leq\rho_{\alpha}(G_{2})\leq n-2<\rho_{\alpha}(K_{2t-1}\vee(K_{n-2t}\cup K_{1})).$
In conclusion, we have
$\rho_{\alpha}(G)\leq\rho_{\alpha}(K_{2t-1}\vee(K_{n-2t}\cup K_{1}))$
with equality if and only if $G=K_{2t-1}\vee(K_{n-2t}\cup K_{1})$. Which is a contradiction to the $A_{\alpha}$-spectral radius condition of Theorem 1.2. This completes the proof of Theorem 1.2. $\Box$
Data availability statement
My manuscript has no associated data.
Declaration of competing interest
The authors declare that they have no conflicts of interest to this work.